OpenStudy (anonymous):

how do you solve log base5 125?

OpenStudy (anonymous):

Well, first I'd look to see if 125 is a power of 5

OpenStudy (angela210793):

\[\log_{5} 125=3\]

OpenStudy (angela210793):

cause 5^3=125

OpenStudy (anonymous):

what aout log\[\log _{2}64 ?\]

OpenStudy (anonymous):

Is 64 a power of 2?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Which power of 2?

OpenStudy (anonymous):

768

OpenStudy (anonymous):

Angela, let her figure it out pls..

OpenStudy (angela210793):

\[\log_{a} b=c \] so that a^c=b

OpenStudy (anonymous):

\[2^{768}=64?\]

OpenStudy (anonymous):

6?

OpenStudy (angela210793):

Wht did I do wrong? why did u deleted it? o.O

OpenStudy (anonymous):

im really not good with this

OpenStudy (angela210793):

delete*

OpenStudy (anonymous):

\[2^6 = 64 \implies log_2(64) = 6\]

OpenStudy (anonymous):

All you have to do is figure out how many 2's you have to multiply to get 64. That number is the log(base 2) of 64.

OpenStudy (anonymous):

i didnt delete it and ok thankyou @polpak & @angela210793

OpenStudy (angela210793):

\[\log_{a} b\]=c cause a^c=b a in this case is 2 and B is 64

OpenStudy (anonymous):

here : log x , base = y :::: x ^ y = base then its log5,125 = 3 ^^

OpenStudy (angela210793):

I know u didn't delete it.... and uw :)

OpenStudy (anonymous):

i think "what is logarithm?" would be a better questiont ... ^^

OpenStudy (angela210793):

yes Korcan agree :)

OpenStudy (angela210793):

http://en.wikipedia.org/wiki/Logarithm

OpenStudy (anonymous):

and now angela help in my questions ?

OpenStudy (angela210793):

still integrals?

OpenStudy (anonymous):

Pleeeaseeeee :) pleeeeaseeee

OpenStudy (anonymous):

I find that this is typically sufficient explanation.. \[log_ba = k \iff b^k = a\]

OpenStudy (angela210793):

@Korcanlet me see if i can solve them @Polpak she didn't understand it that's y i ave her the link...

OpenStudy (anonymous):

polpak help me on my questions plz ? http://openstudy.com/groups/mathematics/updates/4dff4ad80b8bbe4f12e76cf9